OpenStudy (jaydenthequiet):

The areas of two similar rectangles are 80 square feet and 500 square feet. What scale factor applied to the smaller rectangle will give the larger rectangle? A.25/4 B.25/64 C.5/8 D.2/5

OpenStudy (mathstudent55):

Divide the second area by the first area.

OpenStudy (jaydenthequiet):

15/4?

OpenStudy (mathstudent55):

How did you get that?

OpenStudy (mathstudent55):

\(\dfrac{500}{80} = \dfrac{50}{8} = \dfrac{25}{4} \)